OpenStudy (anonymous):

I had this in a lab where HCl+NaOH=H2O+NaCl. The question is at the end of this reaction, how does the (H^+) compare to the (OH)? I don't understand this question, I have the molarity and the volume of the reactants if that helps any. :)

OpenStudy (aaronq):

Was this a titration? do you know the pH of the solution at the end?

OpenStudy (anonymous):

Yes, it was a titration lab but I do not know the pH the molarity of the acid was .2m and the molarity for the base was .1m.

OpenStudy (aaronq):

you mean 0.2 M, right? (make sure you use the right symbol, lowercase "m" means molality) You should first find the pH of the solution, which is a measure of \([H^+]\) Once you do that you can compare the \([OH^-]\) and \([H^+]\), because they're related through Kw. \(K_w=1.0*10^{-14}=[OH^-]*[H^+]\) I think by compare, here, they mean determine the ratio of them, like "there is 2 times as much of one or the other".

OpenStudy (anonymous):

Ok so how would I find the pH of these substances?

OpenStudy (aaronq):

what did you start off with?

OpenStudy (anonymous):

We didn't deal with the pH but we do have 5.3mL of acid and 10.7mL base the acid was HCl and the base was NaOH.

OpenStudy (aaronq):

so you had 5.3mL 0.2 M HCl and 10.7mL 0.1 M NaOH?

OpenStudy (anonymous):

Yes :)

OpenStudy (aaronq):

okay, so to find the pH, we have to use simple stoichiometry. Find the moles of each with the molarity equation: \(M=\dfrac{n_{solute}}{L_{solution}}\) Then cancel them out according to the stoichiometric coefficients, since the ratio is 1:1, there is not much to do. Find the moles.

OpenStudy (anonymous):

Ok so what does the n and L stand for?

OpenStudy (aaronq):

n = moles, L = liters

OpenStudy (anonymous):

Ok got it thank you for the help!!! I 've got to go but I think I can do it from there.

OpenStudy (aaronq):

okay, cool.

OpenStudy (anonymous):

Will you check my answers please?

OpenStudy (anonymous):

@aaronq

OpenStudy (aaronq):

sure thing, it's easier to find a mistake (given you made one) if you post your work as well.

OpenStudy (anonymous):

I'm sure I made one knowing me and chemistry! haha ok so I converted the ml to liters by dividing the mL values by 1000 since 1000 mL per liters. Then I took the two liter values and added them together since you said in the equation the liters of the solution. .0053+.0107=.016 L Then I took that and plugged it into the formula given along with the molarity values: .2=n/.016 and .1=n/.016 then multiply L value to get n by itself for each equation and get Acid: .0032, Base:.0016 I don't know though I think I did it wrong...

OpenStudy (aaronq):

hm theres a problem. When you first find the moles, you're not changing the initial volume, that is you dont add up the volumes just yet. If you do so, it causes the actual number of moles larger than they actually are. You need to first find the moles of each solution.

OpenStudy (anonymous):

Ok I'm sorry I am just so terrible at chemistry and math so stoichiometry is a nightmare for me, I don't even know where to start.

OpenStudy (aaronq):

I know, it can be difficult at first. You just need to understand what is going on, unfortunately it makes time and lost of practice lol So, finding the moles of acid: \(0.2~M=\dfrac{moles}{0.0053~L}\) moles of acid =0.00106 Find the moles of base

OpenStudy (anonymous):

Ok so 0.1=moles/.0107 for base so it's moles= Acid-.00106 and Base- .00107?

OpenStudy (aaronq):

so notice that they are essentially the same. If these cancel out in a 1:1 ratio, how many moles do you have left of each after the cancellations?

OpenStudy (anonymous):

I don't understand...

OpenStudy (aaronq):

if you have 3 friends and each one eats 1 chocolate, and you have 4 chocolates, do you have extra chocolates or hungry friends after you've given all one chocolate?

OpenStudy (anonymous):

You have extra chocolate left.

OpenStudy (aaronq):

so it's the same concept, each mole of acid wants a mole of base, they cancel each other out. So what do you have left after you've cancelled what you can out?

OpenStudy (anonymous):

So I have more base left since .00107 is more than .00106?

OpenStudy (aaronq):

how much exactly?

OpenStudy (anonymous):

1x10^-5

OpenStudy (aaronq):

Thats right. Now that we have the moles of base left, we use the total volume to find the concentration. This is where you need to add up the individual volumes of each. Use the molarity formula again to find the concentration of base with the new volume.

OpenStudy (anonymous):

I'm lost again why would we use that formula and how since we already know the liters of solution, molarity, and moles right?

OpenStudy (aaronq):

what we did first was find the moles of each substance present. When we mixed the two solutions the acid and base reacted to form water and salt ("they cancelled out"). Since we mixed the solutions, the new volume is larger and because we work in concentrations, the volume is important.

OpenStudy (anonymous):

Ok so would I plug in the moles and molarity into the formula to find the new volume?

OpenStudy (aaronq):

nope, we plug in the moles and the volume to find the new molarity.

OpenStudy (anonymous):

Ok so its .00107/.495for base and .00106/.495 for acid?

OpenStudy (aaronq):

hm not quite. remember we "cancelled out" the acid. So all we have left is base, \(1*10^{-5}\) moles of base

OpenStudy (anonymous):

Oh so is it Base: 1.10^-5/.495?

OpenStudy (aaronq):

You have the right idea, but your numbers are wrong. the new volume is 5.3mL+ 10.7mL= 16 mL or 0.016 L so its \(M=\dfrac{1.0*10^{-5}}{0.016~L}=0.000625 =6.25*10^{-4}~M\)

OpenStudy (anonymous):

Ok I see now I just did the adding wrong!

OpenStudy (aaronq):

yep! and to finish the question: This is base right? so it's \([OH^-]=6.25∗10^{−4}~M \)

OpenStudy (aaronq):

use the \(K_w=1.0*10^{-14}=[OH^-]*[H^+]\) to find \([H^+]\)

OpenStudy (anonymous):

Ok thank you!!

OpenStudy (aaronq):

no problem! ask if you any more doubts

OpenStudy (anonymous):

Is it 1.0x10^-14/6.25x10^-4=1.6x10^-19 by any chance?

OpenStudy (anonymous):

@aaronq

OpenStudy (aaronq):

it's \(1.6*10^{-11}\)

OpenStudy (anonymous):

Ok thanks again! Sorry for the slow figuring out of things on my part haha

OpenStudy (aaronq):

no worries! no one is born a genius - at least you're working towards an understanding which will ultimately benefit you in one way or another.

OpenStudy (anonymous):

I hope so haha hopefully in the near future too!

OpenStudy (aaronq):

haha just keep at it!